# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 4.5.2. Main result
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ and $\mathfrak{p}$} a prime of $F$ unramified in $K$. Then we have
@@
This just follows from part (i) of Theorem 4.31 by taking $\chi_{0}=1$ and by choosing an auxiliary prime $\ell\notin S_{\lambda,{\mathfrak{p}}}$ such that $(\ell,2D_{K/F})=1$ and the condition (D) holds.∎
@@ 5. Non-vanishing of Hecke $L$-values modulo $\ell$
This section establishes mod $\ell$ non-vanishing of Hecke $L$-values building on the Rankin–Selberg results in §4. The main results are Theorems 5.8 and 5.9. A bridge among the two non-vanishing arises from a comparison of quaternionic and CM periods (cf. Theorem 5.7), which constitutes the core of this section.
@@ 5.1.1. Setting and strategy
Let $\mathfrak{p}$} be a prime of $F$ unramified in $K$ and $p$ the rational prime below $\mathfrak{p}$}. Recall that $\Gamma_{{\mathfrak{p}}}={\mathrm{Gal}}(K_{\infty}/K)$ is the Galois group of the anticyclotomic ${\mathbb{Z}}_{p}^{[F_{\mathfrak{p}}:{\mathbb{Q}}_{p}]}$-extension of $K$ unramified outside $\mathfrak{p}$}, $\Xi_{\mathfrak{p}}$ the set of $\overline{{\mathbb{Q}}}^{\times}$-valued finite order characters of $\Gamma_{{\mathfrak{p}}}$ and
@@
Let $\ell$ be a prime and fix an embedding $\iota_{\ell}:\overline{{\mathbb{Q}}}\hookrightarrow\overline{{\mathbb{Q}}}_{\ell}$. Suppose that $\ell$ is ordinary for the extension $K/F$, and $\Sigma$ is an $\ell$-ordinary CM type. Then there are CM periods
@@
for any place $v$ of $F$ (cf. Lemma 3.2). Let $\pi_{\lambda}$ be the cuspidal irreducible automorphic representation of $B^{\times}({\mathbb{A}})$ associated to $\phi_{\lambda}$. Let $f_{\lambda}\in\pi_{\lambda}^{\widehat{R}^{\times}}$ be an $\ell$-optimal toric vector defined as in §3.4.
@@ 5.1.2. Lower bound for $\ell$-divisibility of Hecke $L$-values and preliminary mod $\ell$ non-vanishing
For $\ell$ coprime to the conductor of $\lambda$ and $\mathfrak{p}$} split in $K$, this is the content of [26, Prop. 5.1 & Thm. A]. For the remaining cases, see [20, Thm. 3.15, Prop. 4.10 & Thm. 1.5]. ∎
@@
${v}_{\ell}\left(\frac{\pi^{2\kappa}\Gamma_{\Sigma}(\Sigma+\kappa)^{2}\cdot\Omega_{\lambda}}{D_{F}\Omega_{\infty}^{2\Sigma+4\kappa}}\prod_{w\in\Sigma_{\ell}}G(\lambda_{w})^{2}\right)\geq 2\sum_{\begin{subarray}{c}v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))\text{ inert}\end{subarray}}\mu_{\ell}(\lambda_{v}).$
@@
${v}_{\ell}\left(\frac{\sqrt{|D_{K}|}\cdot L(1/2,\pi_{\lambda,K}\otimes\nu)}{\Omega_{\lambda}}\right)={v}_{\ell}\left(\frac{\sqrt{|D_{K}|}\cdot L(1,\lambda\nu)L(1,\lambda\nu^{-1})}{\Omega_{\lambda}}\right)=0$
@@
Hence, the result follows. ∎
In the following we consider a lower bound for the $\ell$-adic valuation of Hecke $L$-values in the complementary case that $\deg{\mathfrak{p}}>1$.
Given a positive integer $a$, recall that we have a partition of $\Xi_{{\mathfrak{p}}}$ into finitely many subfamilies
$\Xi_{\mathfrak{p}}\doteq\bigsqcup_{\mu:O_{{\mathfrak{p}}}/p^{a}\twoheadrightarrow{\mathbb{Z}}/p^{a}{\mathbb{Z}}}\Xi_{{\mathfrak{p}}}(\mu)$
(cf. §4.5.2). Employing Hida’s approach, one is led to a refinement of Theorem 5.2 under the following condition (A):
[…]
$\bullet$ $\sum_{\begin{subarray}{c}v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))\text{ inert}\\ v\nmid{\mathfrak{p}}\end{subarray}}\mu_{\ell}(\lambda_{v})=0.$
The main result of this subsection is the following.
Theorem 5.4.
Let $K$ be a CM field, $F$ its maximal totally real subfield and $\Sigma$ a CM type. Let $\lambda$ be a self-dual Hecke character over $K$ of infinity type $\Sigma+\kappa(1-c)$ for $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$. Let $\ell\nmid 2D_{F}$ be a prime such that $\Sigma$ is an $\ell$-ordinary CM type. Let $\mathfrak{p}$} be a prime of $F$ such that ${\mathfrak{p}}\nmid\ell D_{K/F}$. Suppose that the condition (A) holds.
  • i)
    Then we have
    for all except finitely many $\nu\in\Xi_{\lambda,{\mathfrak{p}}}^{+}$.
  • ii)
    Fix an integer $a$ such that ${\bf k}_{\ell}^{\times}[p^{\infty}]=\mu_{p^{a}}$ as in (4.12) and let $\mu:O_{{\mathfrak{p}}}/{\mathfrak{p}}^{ae}\twoheadrightarrow\mu_{p^{a}}$ be a character such that $\Xi_{\lambda,{\mathfrak{p}}}^{+}\cap\Xi_{{\mathfrak{p}}}(\mu)=\infty$. Then there exists a Zariski dense subset of $\Xi_{\lambda,{\mathfrak{p}}}^{+}\cap\Xi_{{\mathfrak{p}}}(\mu)$ consisting of characters $\nu$ such that
Proof.
The result basically follows from [26] and [20] (cf. [21, 25]).
As for part (i), Suppose first that ${\mathfrak{p}}\mathrm{cond}(\lambda)$ is divisible by a prime of $F$ split in $K$. Then there is an $\ell$-adic Hilbert Eisenstein series $\widehat{{\mathbb{E}}}_{\lambda,\Sigma}$ whose constant term at a cusp is 0, and such that the associated $\ell$-integral measure $\mathcal{E}$} on $\Xi_{{\mathfrak{p}}}$ interpolates [^21] This follows by taking $S_{\text{split},2}$ in loc. cit. to be the empty set.
$L^{{\mathrm{alg}}}(1,\lambda\nu)$
for all except finitely many $\nu\in\Xi_{{\mathfrak{p}}}$ (cf. [20, Prop. 3.13]).
In view of the $q$-expansion principle (cf. [20, Prop. 2.6]) and the lower bound for $\ell$-adic valuation of non-constant Fourier coefficients of $\widehat{{\mathbb{E}}}_{\lambda,\Sigma}$ as in  [20, Prop. 4.10], the $\ell$-adic valuation of the measure $\mathcal{E}$} is bounded below by
$\sum_{\begin{subarray}{c}v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))\text{ inert}\\ v\nmid{\mathfrak{p}}\end{subarray}}\mu_{\ell}(\lambda_{v}),$
yielding part (i) in this case.
Next suppose that $(\ell,2D_{F})=1$ and ${\mathfrak{p}}\mathrm{cond}(\lambda)$ is not divisible by a prime of $F$ split in $K$. Then the Eisenstein series ${\mathbb{E}}_{\lambda,\Sigma}$ is still $\ell$-integral by the $q$-expansion principle due to $\ell$-integrality [^22] This follows by taking $S_{\text{split},2}$ in the definition of ${\mathbb{E}}_{\lambda,\Sigma}$ in Section 3.6 of [20] to be the empty set. of non-constant terms (cf. [20, Prop. 4.10]) and the fact that the $\ell$-adic valuation of constant terms is at least ${v}_{\ell}(L(0,\eta_{K/F}))$ (cf. [26, Prop. 4.7]). Thus the interpolation as in [20, Prop. 3.13] still holds. In turn, the aforementioned argument applies for almost all $\nu$, yielding $\ell$-integrality of the measure $\mathcal{E}$}.
As for part (ii), the argument in Section 5 of [20] applies for the subfamily $\Xi(\mu)$, leading to part (ii).
@@
  • ii)
    In this paper we only use part i) of Theorem 5.4 pertaining to the case $\deg{\mathfrak{p}}=1$ for our general non-vanishing results, i.e. for the case $\deg{\mathfrak{p}}=1$ and indirectly for the case $\deg{\mathfrak{p}}>1$.
@@ 5.2. Comparison of periods and $(\ell,p)$ non-vanishing
  • b’)
    Let ${\mathfrak{p}}\nmid\ell\mathfrak{N}_{\lambda}^{-}$ be a prime of $F$ of degree at least $2$. Fix an integer $a$ such that ${\bf k}_{\ell}^{\times}[p^{\infty}]=\mu_{p^{a}}$ as in (4.12). Let $\mu:O_{{\mathfrak{p}}}/{\mathfrak{p}}^{ae}\twoheadrightarrow\mu_{p^{a}}$ be a character with $\Xi_{\lambda,{\mathfrak{p}}}^{+}\cap\Xi_{{\mathfrak{p}}}(\mu)=\infty$ and $\delta_{{\mathfrak{p}},\mu}$ be as in Definition 4.26. Then
    ${v}_{\ell}\left(L^{{\mathrm{alg}}}(1,\lambda\nu)L^{{\mathrm{alg}}}(1,\lambda\nu^{-1})\right)={v}_{\ell}(\delta_{{\mathfrak{p}},\mu})+2\sum_{\text{$v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))$ inert}}\mu_{\ell}(\lambda_{v})$
    for all except finitely many $\nu\in\Xi_{\lambda,{\mathfrak{p}}}^{+}\cap\Xi_{{\mathfrak{p}}}(\mu)$.
@@
We first prove that a) $\Leftrightarrow$ b). Let $\mathfrak{q}$} be as in part b).
Based on the mod $\ell$ non-vanishing of Rankin–Selberg $L$-value as in Theorems 4.14 and 4.17 for the self-dual pair $(\pi_{\lambda},\chi_{0}\nu)$, we have
@@
where we have used the fact that ${v}_{\ell}(\prod_{w\in\Sigma_{\ell}}G(\lambda_{w}\chi_{0,w}))={v}_{\ell}(\prod_{w\in\Sigma_{\ell}}G(\lambda_{w}))$ since $\chi_{0}$ is unramified at $\ell$, and $\mu_{\ell}(\lambda_{v}\chi_{0,v})=\mu_{\ell}(\lambda_{v})$ for $v|N_{K/F}({\mathrm{Cond}}(\lambda))$ inert in $K$. Hence, in light of (5.2), (5.3) and (5.4), part b) follows.
By the same argument, the equivalence of parts a) and b’) is a consequence of the non-vanishing of Rankin–Selberg $L$-values as in Theorem 4.31, where the invariant $\delta_{{\mathfrak{p}},\mu}$ appears. ∎
@@
Such a $\chi_{0}$ exists since $\mathfrak{q}$} is inert in $K$ by the epsilon factor formula (cf. Lemma 4.4.1). If $\varepsilon(\lambda)=+1$, then we simply take $\chi_{0}=1$ and if $\varepsilon(\lambda)=-1$, we take $\chi_{0}$ to be sufficiently ramified.
Choose an auxiliary degree one prime $\mathfrak{r}$} of $F$ such that
@@
    the rational prime $r$ below it is coprime to $q\ell{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$ and has odd residue characteristic.
Since $\varepsilon(\lambda\chi_{0})=\varepsilon(\lambda\chi_{0}^{-1})=+1$, applying the mod $\ell$ non-vanishing as in part (ii) of Theorem 5.2 for $\lambda\chi_{0}$ and $\lambda\chi_{0}^{-1}$, there exist characters $\nu_{1},\nu_{2}\in\Xi_{{\mathfrak{r}}}$ such that
@@
Since $\mathfrak{r}$} has odd residual characteristic, one can find [^24] That is, $\nu_{0}=\sqrt{\nu_{1}}\sqrt{\nu_{2}}$ and $\nu=\sqrt{\nu_{1}}\sqrt{\nu_{2}}^{-1}$, where $\sqrt{\nu_{i}}$ is a square-root of $\nu_{i}$ in $\Xi_{{\mathfrak{r}}}$. $\nu_{0},\nu\in\Xi_{{\mathfrak{r}}}$ such that $\nu_{0}\nu=\nu_{1}$ and $\nu_{0}{\nu}^{-1}=\nu_{2}$. Then we have
@@
i.e. part (a) of Proposition 5.6 for the Hecke character $\lambda\nu_{0}$. Consequently, a series of implications ensues: part (b) of Proposition 5.6 for Hecke characters $\lambda\nu_{0}\chi_{0}$ and $\lambda\nu_{0}\chi_{0}^{-1}$ holds by the same proposition (prime $\mathfrak{p}$} therein being $\mathfrak{r}$} and $\chi_{0},{\mathfrak{q}}$ being as above). Then the latter implies part (b) of Proposition 5.6 for Hecke characters $\lambda\chi_{0}$ and $\lambda\chi_{0}^{-1}$ (prime $\mathfrak{p}$} therein being $\mathfrak{r}$} and $\chi_{0},{\mathfrak{q}}$ as above). Hence, part (a) of Proposition 5.6 for the Hecke character $\lambda$ follows and so
@@
Since $\mathfrak{r}$} splits in $K$, we can further choose Hecke characters $\nu_{1}$ and $\nu_{2}$ such that
@@
    ${\mathrm{ord}}_{{\mathfrak{r}}}\mathrm{cond}(\nu)\geq{\mathrm{ord}}_{\mathfrak{r}}\mathrm{cond}(\pi_{\pi\nu_{0}})$ and,
@@
for all but finitely many $\mu\in\Xi_{{\mathfrak{r}}}$. Together with the lower bound for the $\ell$-adic valuation of Hecke $L$-values as in part (i) of Theorem 5.2, we then have [^25] Here we use the fact that an unramified finite order twist at $\ell$ does not change the $\ell$-adic valuation of the underlying Gauss sum.
@@
5.3. Non-vanishing results
5.3.1. The case $\deg{{\mathfrak{p}}}=1$

### v2

@@ 4.4.4. Consequence on non-vanishing of Hecke L-values
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ and $\mathfrak{p}$} a degree one prime of $F$ unramified in $K$. Then we have
@@
This just follows from Theorem 4.15 and Theorem 4.19 for $\mathfrak{p}$} inert and split in $K/F$ respectively, by taking $\chi_{0}=1$ and by choosing an auxiliary prime $\ell$ such that $(\ell,2{\mathfrak{p}}D_{K/F})=1$.∎
@@ 5. Non-vanishing of Hecke $L$-values modulo $\ell$
This section establishes mod $\ell$ non-vanishing of Hecke $L$-values building on the Rankin–Selberg results in §4. The main result is Theorem 5.7. A bridge among the two non-vanishing arises from a comparison of quaternionic and CM periods (cf. Theorem 5.6), which constitutes the core of this section.
@@ 5.1.1. Setting and strategy
Let $\mathfrak{p}$} be a prime of $F$ and $p$ the rational prime below $\mathfrak{p}$}. Recall that $\Gamma_{{\mathfrak{p}}}={\mathrm{Gal}}(K_{\infty}/K)$ is the Galois group of the anticyclotomic ${\mathbb{Z}}_{p}^{[F_{\mathfrak{p}}:{\mathbb{Q}}_{p}]}$-extension of $K$ unramified outside $\mathfrak{p}$}, $\Xi_{\mathfrak{p}}$ the set of $\overline{{\mathbb{Q}}}^{\times}$-valued finite order characters of $\Gamma_{{\mathfrak{p}}}$ and
@@
Fix an embedding $\iota_{\infty}:\overline{{\mathbb{Q}}}\hookrightarrow{\mathbb{C}}$. Let $\ell$ be a prime and fix an embedding $\iota_{\ell}:\overline{{\mathbb{Q}}}\hookrightarrow\overline{{\mathbb{Q}}}_{\ell}$. Suppose that $\ell$ is ordinary for the extension $K/F$, and $\Sigma$ is an $\ell$-ordinary CM type. Then there are CM periods
@@
For the rest of this section, we will assume that
  • $\mathfrak{p}$} is a degree one prime of $F$ that unramified in $K/F$,
  • $\ell\nmid 2{\mathfrak{p}}D_{F}$ is ordinary.
@@
for any place $v$ of $F$ (cf. Lemma 3.2). Let $\pi_{\lambda}$ be the cuspidal irreducible automorphic representation of $B^{\times}({\mathbb{A}})$ associated to $\phi_{\lambda}$. Let $f_{\lambda}\in\pi_{\lambda}^{\widehat{R}^{\times}}$ be an $\ell$-optimal toric vector defined as in §3.4, and we take $S_{0}$ as in §3.4 to be coprime to $\mathfrak{N}_{\lambda}^{-}$. Then $\phi_{\lambda}$ does not depends on $S_{0}$.
@@ 5.1.2. Lower bound for $\ell$-divisibility of Hecke $L$-values and preliminary mod $\ell$ non-vanishing
For $\ell$ coprime to the conductor of $\lambda$ and $\mathfrak{p}$} split in $K$, this is the content of [25, Prop. 5.1 & Thm. A]. For the remaining cases, see [19, Thm. 3.15, Prop. 4.10 & Thm. 1.5]. See also the remark at the beginning of Section 6.1 of [25] for more explanation about part i).
@@
${v}_{\ell}\left(\frac{\pi^{2\kappa}\Gamma_{\Sigma}(\Sigma+\kappa)^{2}\cdot\Omega_{\lambda}}{\Omega_{\infty}^{2\Sigma+4\kappa}}\prod_{w\in\Sigma_{\ell}}G(\lambda_{w})^{2}\right)\geq 2\sum_{\begin{subarray}{c}v|{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))\text{ inert}\end{subarray}}\mu_{\ell}(\lambda_{v}).$
@@
${v}_{\ell}\left(\frac{\sqrt{|D_{K}|}\cdot L(1,\lambda\nu)L(1,\lambda\nu^{-1})}{\Omega_{\lambda}}\right)={v}_{\ell}\left(\frac{\sqrt{|D_{K}|}\cdot L(1/2,\pi_{\lambda,K}\otimes\nu)}{\Omega_{\lambda}}\right)=0$
@@
Note that the assumption on $\ell$ implies that $D_{K}$ is an $\ell$-adic unit. Hence, the result follows. ∎
@@ 5.2. Comparison of periods and $(\ell,p)$ non-vanishing
Let $\mathfrak{q}$} be as in part b).
Based on the mod $\ell$ non-vanishing of Rankin–Selberg $L$-value as in Theorems 4.15 and 4.19 for the self-dual pair $(\pi_{\lambda},\chi_{0}\nu)$, we have
@@
where we have used the fact that ${v}_{\ell}(\prod_{w\in\Sigma_{\ell}}G(\lambda_{w}\chi_{0,w}))={v}_{\ell}(\prod_{w\in\Sigma_{\ell}}G(\lambda_{w}))$ since $\chi_{0}$ is unramified at $\ell$, and $\mu_{\ell}(\lambda_{v}\chi_{0,v})=\mu_{\ell}(\lambda_{v})$ for $v|{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$ inert in $K$. Hence, in light of (5.2), (5.3) and (5.4), part b) follows.
@@
Such a $\chi_{0}$ exists since $\mathfrak{q}$} is inert in $K$ by the epsilon factor formula (cf. Lemma 4.4.1). If $\varepsilon(\lambda)=+1$, then we simply take $\chi_{0}=1$ and if $\varepsilon(\lambda)=-1$, we take $\chi_{0}$ to be sufficiently ramified. Note that the assumption $\ell\neq 3$ implies the existence of a prime $\mathfrak{q}$} with $\ell\nmid q^{2}-1$.
Choose an auxiliary degree one prime $\mathfrak{q}$$^{\prime}$ of $F$ such that
@@
    the rational prime $q^{\prime}$ below it is coprime to $q\ell{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$ and has odd residue characteristic.
Since $\varepsilon(\lambda\chi_{0})=\varepsilon(\lambda\chi_{0}^{-1})=+1$, applying the mod $\ell$ non-vanishing as in part (ii) of Theorem 5.2 for $\lambda\chi_{0}$ and $\lambda\chi_{0}^{-1}$, there exist characters $\nu_{1},\nu_{2}\in\Xi_{{\mathfrak{q}}^{\prime}}$ such that
@@
Since $\mathfrak{q}$$^{\prime}$ has odd residual characteristic, one can find [^24] That is, $\nu_{0}=\sqrt{\nu_{1}}\sqrt{\nu_{2}}$ and $\nu=\sqrt{\nu_{1}}\sqrt{\nu_{2}}^{-1}$, where $\sqrt{\nu_{i}}$ is a square-root of $\nu_{i}$ in $\Xi_{{\mathfrak{q}}^{\prime}}$. $\nu_{0},\nu\in\Xi_{{\mathfrak{q}}^{\prime}}$ such that $\nu_{0}\nu=\nu_{1}$ and $\nu_{0}{\nu}^{-1}=\nu_{2}$. Then we have
@@
i.e. part (a) of Proposition 5.5 for the Hecke character $\lambda\nu_{0}$. Consequently, a series of implications ensues: part (b) of Proposition 5.5 for Hecke characters $\lambda\nu_{0}\chi_{0}$ and $\lambda\nu_{0}\chi_{0}^{-1}$ holds by the same proposition (prime $\mathfrak{p}$} therein being $\mathfrak{q}$$^{\prime}$ and $\chi_{0},{\mathfrak{q}}$ being as above). Then the latter implies part (b) of Proposition 5.5 for Hecke characters $\lambda\chi_{0}$ and $\lambda\chi_{0}^{-1}$ (prime $\mathfrak{p}$} therein being $\mathfrak{q}$$^{\prime}$ and $\chi_{0},{\mathfrak{q}}$ as above). Hence, part (a) of Proposition 5.5 for the Hecke character $\lambda$ follows and so
@@
Since $\mathfrak{q}$$^{\prime}$ splits in $K$, we can further choose Hecke characters $\nu_{1}$ and $\nu_{2}$ such that
@@
    ${\mathrm{ord}}_{{\mathfrak{q}}^{\prime}}\mathrm{cond}(\nu)\geq{\mathrm{ord}}_{\mathfrak{q}}^{\prime}\mathrm{cond}(\pi_{\pi\nu_{0}})$ and,
@@
for all but finitely many $\mu\in\Xi_{{\mathfrak{q}}^{\prime}}$. Together with the lower bound for the $\ell$-adic valuation of Hecke $L$-values as in part (i) of Theorem 5.2, we then have [^25] Here we use the fact that an unramified finite order twist at $\ell$ does not change the $\ell$-adic valuation of the underlying Gauss sum.
@@
5.3. Mod $\ell$ non-vanishing
The main result of this section is the following.